1. A stock sells at $15 per share. a.) What is the EPS for the company if it has a P/E ratio of 20?

b.) If the company's dividend yield is 3%, what is its dividend per share?

c.) What is the book value of the company if the price-to-book ratio is 1.5 and it has 100,000 shares of stock outstanding.
Answer #1

a)

current price per share , p = $15

(P/E) = 20

EPS = p/(P/E) = 15/20 = $0.75

b)

dividend yield , d = 3% = 0.03

dividend per share = d*p = 0.03*15 = $0.45

c)

price to book ratio , r = 1.5

value of shares = v = p*100,000 shares = 15*100,000 = 1,500,000

Book value of the company = v/ r = 1,500,000/1.5 = 1,000,000